### v1

Abstract.
This paper presents the first rotating leader state machine replication (SMR) protocol that allows transactions to be confirmed in just a single round-trip time in the Byzantine fault tolerance (BFT) setting. Based on minimal alterations to the Internet Computer Consensus (ICC) protocol and with negligible communication overhead, we introduce a novel dual mode mechanism that enables optimal block finalization latency in the fast path. Crucially, the modes of operation are integrated, such that even if the fast path is not effective, no penalties are incurred. Moreover, our algorithm maintains the core attributes of the original ICC protocol, including optimistic responsiveness and rotating leaders without the necessity for a view-change protocol.
We prove the correctness of our Fast Internet Computer Consensus (FICC) protocol and provide an open-source implementation of it. Both the FICC and original ICC protocol are compared in a globally distributed wide-area network. Our evaluation reveals that the FICC protocol achieves reduced latency compared to the ICC protocol, without requiring additional security assumptions. Furthermore, by increasing the number of replicas to $n=5f+1$, we exhibit that latency improvements close to the theoretical maximum of 33% are attainable. We conclude by highlighting the network topology as a significant factor in evaluating and comparing the latency of consensus algorithms.
state machine replication, consensus, byzantine fault tolerance, blockchain, fast path
copyright: nonejournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYccs: Security and privacy Distributed systems securityccs: Applied computing Secure online transactions
1. Introduction
Byzantine fault tolerance (BFT) is a class of protocols that provide guarantees in the presence of arbitrary faults, such as a powerful adversary controlling both a share of participants and the network scheduling (Lamport et al., 1982). BFT protocols can be applied whenever a fixed (also called permissioned) set of untrusted parties desires to reach an agreement.
Popularized by Bitcoin (Nakamoto, 2009), Ethereum (Buterin, 2013) and other digital currencies, protocols that tolerate an open set of participants have emerged. This has not lessened interest in the permissioned setting, however, as their approaches are often based on advancements from permissioned counterparts (see Ethereum’s Casper (Buterin and Griffith, 2017) and Algorand (Gilad et al., 2017) for instance).
The practicality of these systems has been shown in the context of finance and digital art (Uniswap, 2023; Compound, 2023; MakerDAO, 2023; OpenSea, 2023), but other promising applications such as file storage, social media, governance, and democracy have emerged (Filecoin, 2023; dscvr, 2023; Tally, 2023; Worldcoin, 2023). In essence, BFT protocols can be applied whenever multiple parties desire to perform some computation, without trusting any set of participants fully.
These decentralized systems are not unlike a decentralized world computer, enabling anyone to run Turing-complete programs. Various such world computers exist today (Buterin, 2013; Dfinity, 2023c; Filecoin, 2023; Blackshear et al., 2023), with varying trade-offs regarding performance, security, and inclusiveness. At each system’s core, lays an ordering protocol assuring that resource accesses are consistent across all participants (called replicas), thus guaranteeing deterministic and secure execution of the computation.
Byzantine agreement (BA) protocols typically provide consensus on a single decision, while state machine replication protocols (SMR) focus on making efficient use of this costly primitive. In practice, the most widely used SMR protocols (such as Bitcoin (Nakamoto, 2009), Ethereum (Buterin, 2013), and Algorand (Gilad et al., 2017)) make use of an elected leader to propose a batch of transactions, called a block. A total order across leaders is then obtained by chaining blocks.
There are a few considerations that are especially crucial to such leader-based protocols. Firstly, leaders can misbehave and propose conflicting blocks. While this can be detected, the resulting change of the leader (called view-change in the literature) leads to notoriously high complexity. Honest, but slow leaders are just as problematic, as they cannot easily be called out, but can stall the effective throughput nonetheless (Clement et al., 2009).
Moreover, relying on a single leader leads to uneven load, both in terms of computation and communication (Crain et al., 2018). Finally, leaders can sometimes extract value from their control of the block creation. On the Ethereum blockchain for example, miner-extractable value (MEV) is a major source of income for replicas and leads to potentially dangerous incentives to deviate from honest behavior (Daian et al., 2020).
These issues can at least partially be resolved by employing rotating or random leaders (Camenisch et al., 2022), which is the class of protocols we aim to improve upon in this work.
Figure 1. The FICC protocol can terminate after just two communication steps. Existing rotating leader SMR protocols require at least three communication steps.
Specifically, we address what we believe to be a major roadblock in the adoption of decentralized computers: high latency. In contrast to privacy and security, latency is at the forefront of the user experience. While centralized and trusted applications (such as credit card payments or social media) provide almost instantaneous services, their Byzantine fault tolerant counterpart typically comes at the cost of a noticeable delay, in some cases in the order of minutes or even hours (such as in Bitcoin). Furthermore, contrary to other metrics such as throughput, latency cannot be readily improved by hardware upgrades.
In this work, we present the first rotating leader SMR protocol that allows transactions to be confirmed in just a single round-trip time. This is optimal, as to achieve fault tolerance the block must 1) reach a fraction of replicas, that have to 2) respond to acquiesce to the processing. This is visualized in fig. 1.
We are able to achieve this by uniting two lines of research. The first deals with the theoretical bounds of latency, and has established optimally fast single-shot protocols in the good case (Kursawe, 2002; Martin and Alvisi, 2006; Aublin et al., 2015; Kuznetsov et al., 2021; Abraham et al., 2021b). The second, more well-known line of research, provides a class of consensus protocols that solve state machine replication, in an efficient and fair way by employing rotating leaders. Prominent examples include HotStuff, Tendermint, Casper, Algorand and Bullshark (Yin et al., 2019; Buchman, 2016; Buterin and Griffith, 2017; Gilad et al., 2017; Spiegelman et al., 2022).
As Chan and Pass point out (Chan and Pass, 2023), rotating leader protocols can be differentiated by either favoring proposal confirmation time, meaning block finalization latency, or block creation time, corresponding to the delay between the blocks and thus the associated throughput. The Internet Computer Consensus family of protocols (Camenisch et al., 2022) excels in both of these metrics. The protocol powers the Internet Computer (IC) project (Dfinity, 2023c): a decentralized platform enabling developers to deploy dApps that run fully on the blockchain. The system is composed of multiple permissioned sets of replicas running datacenter-grade infrastructure managed by a decentralized autonomous organization.
Exhibiting around one to two seconds of block finalization latency, this deployment of ICC is among the quickest consensus algorithms in the blockchain space. Yet, it is arguably not fast enough to compete with the Web services hosted on centralized server infrastructures.
Our contribution. We present the Fast Internet Computer Consensus (FICC) protocol that achieves optimal one-shot consensus latency while preserving the state-of-the-art characteristics of ICC.
  1. (1)
    FICC is the first SMR protocol with rotating leaders supporting two-step chain growth and two-step finalization under partial synchrony. We prove the safety and liveness of FICC for $n\geq 3f+2p+1$, where $n$ is the total number of replicas in the network, and $f$ is the maximum number of Byzantine replicas tolerated. Additionally, $p$ is a parameter that can be set freely ($0\leq p\leq f$), and determines the effectiveness of the fast path. We prove that as long as no more than $p$ replicas are unresponsive, FICC block latency is just two network delays.
  2. (2)
    We provide an experimental FICC implementation in Rust, forked from the Internet Computer codebase.
  3. (3)
    We measure the proposal finalization time in a wide area network spanning the globe. We empirically show that FICC improves proposal finalization time compared to the reference ICC implementation without any relaxation of the security guarantees.
2. Related work
Byzantine fault tolerance was introduced by Lamport et al. in 1982 (Lamport et al., 1982). A long line of work studies consensus protocols in the permissioned setting under various synchrony assumptions (Ben-Or, 1983; Bracha, 1987; Lamport, 2003)). The partially synchronous network model was first introduced in (Dwork et al., 1988), together with the first consensus protocol for this setting. The first truly practical protocol in the partially asynchronous setting is the PBFT protocol, as described in (Castro et al., 1999).
Fast Consensus. Our algorithm relies on a long line of established work on reaching consensus in two communication steps (or one round trip time), also referred to as fast or early-stopping consensus (Brasileiro et al., 2001; Kursawe, 2002; Song and van Renesse, 2008; Martin and Alvisi, 2006; Guerraoui and Vukolić, 2007). As shown by Brasileiro et al. (Brasileiro et al., 2001) in the crash-fault model, it is possible to terminate early if “enough” acknowledgments are received. A year later, a fast path was presented by Kursawe (Kursawe, 2002), in which a two-step fast path is paired with a subprotocol in the slow path. Upon the expiration of a timer, a fallback subprotocol is used to ensure liveness.
Song et al. describe Bosco (Song and van Renesse, 2008), an algorithm that provides a two-step fast path to any underlying consensus algorithm. Their fast path does not rely on synchrony assumptions, and, assuming $n>5f$, is triggered when all servers are in pre-agreement. Assuming $n>7f$, the fast path is triggered when all honest servers are in pre-agreement instead. Using similar building blocks it was later shown that consensus can be performed “on demand”, i.e., the fallback is only used when the fast path does not succeed (Sliwinski et al., 2022). As fig. 2 visualizes, these approaches come at the cost of a longer recovery during the fallback on the slow path though, as the slow path can only be started once the fast path round is over. In this work, we show that both the (single) fast path round and the first slow path round can happen simultaneously (with only constant message overhead).
Figure 2. The different approaches of fast path protocols. In many protocols, the slow path starts after the fast path fails. In FICC, the fast path is integrated into the slow path.
Earlier, Martin and Alvisi introduce Fast Byzantine Paxos (FaB Paxos) (Martin and Alvisi, 2006), a family of protocols parameterized not only by $f$ but also $p$ ($0\leq p\leq f$), the number of (non-leader) replicas that are not needed for the fast path to succeed. Their algorithms provide Byzantine agreement, when $n\geq 3f+2p+1$, where the fast path triggers as long as $n-p$ servers behave honestly and are in pre-agreement. We use the same security model and guarantees in this work, however, we do not require progress certificates to guarantee safety. Additionally, instead of relying on a variant of PBFT in the slow path, we integrate our protocol with ICC, sidestepping the complexity introduced by view changes.
The lower bound by Martin and Alvisi (Martin and Alvisi, 2006) has recently been improved to $n\geq max(3f+2p-1,3f+1)$ by Kuznetsov et al. (Kuznetsov et al., 2021) and Abraham et al. (Abraham et al., 2021b). This is applicable when servers and clients are identical sets. Their insight to observe and discard actions by provably misbehaving leaders does not trivially apply to our use case, as we propose an SMR algorithm instead of a single-shot consensus instance. A more rigorous analysis is left for future work.
Fast SMR. Ideas from the FaB protocol were borrowed by the Zyzzyva protocol (Kotla et al., 2007) as well, which turns the single-shot FaB consensus into an SMR protocol. Improving upon Zyzzyva, Aliph (Aublin et al., 2015) reduces the latency to two steps. These protocols do not allow running both the fast path and a slow path concurrently though, and thus suffer from switching costs (see fig. 2). Zyzzyva safety issues were pointed out by (Abraham et al., 2017) and lead to the development of Zelma (Abraham et al., 2018), combining FaB and Zyzzyva’s benefits.
Zelma is also at the core of the SBFT protocol (Gueta et al., 2019). SBFT is the first protocol running both pessimistic and optimistic paths simultaneously in a dual mode. However, it requires more communication steps when compared to FICC. Indeed, this approach lends itself to ICC exceptionally well, as the protocol tolerates arbitrarily delaying the decision on a value until (in a phase of synchrony) a successful consensus instance retroactively finalizes previous instances. A large body of work has based itself on the 3-phase (6-step) HotStuff (Yin et al., 2019) protocol and presented improvements upon it. Jolteon and Ditto (Gelashvili et al., 2022) have improved the latency while adding an asynchronous fallback to enhance its performance during epoch synchronization (see table 1).
Most recently Malkhi and Nayak propose a 2-phase (4-step) protocol and a simpler fast path (Malkhi and Nayak, 2023). It is our understanding that these protocols reach no lower than $3\delta$ proposer latency in the fast path.
Rotating Leader BFT protocols. Mir-BFT (Stathakopoulou et al., 2019) and its follow-up work (Stathakopoulou et al., 2022) improve upon sequential-leader approaches, by running PBFT instances on a set of leaders. When a leader is slow, it is replaced. Compared to (F)ICC protocols, it suffers from the complexity of view-changes. Other SMR protocols are designed to perform well under attack (Avarikioti et al., 2023). However, in the permissioned blockchain setting it is often assumed that during periods of synchrony, misbehavior can be punished, either by exclusion or slashing (Buterin and Griffith, 2017). As such we focus on the latency in the optimistic case and satisfy the same pessimistic liveness as ICC, which was shown to be adequate in (Chan and Pass, 2023). Recent concurrent work by Chan and Pass (Chan and Pass, 2023) provides a theoretical framework for this class of algorithms, and table 1 is partially inspired by it. Their proposed protocol called Simplex falls into the same category as the ICC and FICC algorithms but improves pessimistic liveness by only allowing a single leader per round. We believe that a technique similar to the one shown in this work could be applied to bring our fast path to Simplex too. In (Abraham et al., 2021a) Abraham et al. prove that $2\Delta$ is the upper and lower bound for the good-case latency of rotating leader BFT protocols, in the synchronous setting. In this work, we provide a matching upper bound in the partially synchronous setting.
DAG-based BFT protocols. Another line of work proposes to solve consensus efficiently by using DAG-based protocols (Keidar et al., 2021; Spiegelman et al., 2022; Babel et al., 2023; Malkhi et al., 2023).
One main idea is to disconnect transactions broadcasting from finalization (Danezis et al., 2022), and another is to allow the finalization of blocks outside the main blockchain by causally referencing them (Lewenberg et al., 2015; Keidar et al., 2021). In Bullshark (Spiegelman et al., 2022), one of the most recent protocols, all replicas are parallel leaders and broadcast one batch of transactions in each round using Byzantine Consistent Broadcast. While doing so, they include references to at least $n-f$ blocks from the previous round. A more recent protocol called BBCA-Chain (Malkhi et al., 2023) reduces Bullshark’s latency by removing the need for specialized block layers, and instead broadcasts blocks using a primitive called Byzantine Broadcast with Complete-Adopt (BBCA). (F)ICC does not allow causally referencing non-leader blocks to finalize them and thus doesn’t benefit from the same throughput advantages. However, we believe ICC is a great fit to describe our fast path mechanism, which might be applicable to other DAG protocols too.
In table 1 an overview of state-of-the-art protocols is shown. For both the proposer latency and the block creation latency, we have included the required number of replicas that have to respond in order to proceed in the optimistic case (synchronous round with a block proposed by an honest leader). While irrelevant in the calculation of the theoretical latency bounds, we found that small changes in these requirements could have large effects in practice, especially on global-scale deployments as they are seen today. (Intuitively, it is not uncommon for a few outlier nodes to have a higher latency. If progress can be made safely without them, a large performance increase can be experienced.)
Block Finalization
Latency
Finalization
Requirement
Block Creation
Latency
Block Creation
Requirement
Number of
Replicas $\geq$
Supports
Rotating
Leaders
Casper FFG (Buterin and Griffith, 2017) $O(\Delta)$ $\ {{\lceil\frac{2n}{3}\rceil}}$ $O(\Delta)$ N/A *
$13f+$ 1
Jolteon (Gelashvili et al., 2022)
Pipelined
Fast HotStuff (Jalalzai et al., 2020)
$5\delta$ $12f+$ 1 $2\delta$ $12f+$ 1 $13f+$ 1
PaLa (Chan et al., 2018) $4\delta$ ${\lceil\frac{2n}{3}\rceil}$ $2\delta$ ${\lceil\frac{2n}{3}\rceil}$ $13f+$ 1
Zelma (Abraham et al., 2018) $2\delta$ $n-p$ $2\delta$ $n-f-p$ $13f+2p+$ 1
SBFT (Gueta et al., 2019) $3\delta$ $13f+p+$ 1 $3\delta$ $13f+p+$ 1 $13f+2p+$ 1
Bullshark (Spiegelman et al., 2022) $4\delta$
$n-f$ with
$f+1$ overlap
$2\delta$
$n-f$ (odd rounds)
$f+1$ / $12f+$ 1 (even rounds)
$13f+$ 1
BBCA-Chain (Malkhi et al., 2023) $3\delta$ $12f+$ 1 $3\delta$ $12f+$ 1 $13f+$ 1
ICC (Camenisch et al., 2022)
Simplex (Chan and Pass, 2023)
$3\delta$ $n-f$ $2\delta$ $n-f$ $13f+$ 1
FICC $2\delta$ $n-p$ $2\delta$ ${\lfloor\frac{n+f}{2}\rfloor}+1$ $13f+2p+$ 1
Table 1. Popular and Fast State Machine Protocols. *Non-equivocation is enforced by slashing. † To the best of our knowledge no pipelining is specified. $\Delta$ denotes the message delivery time upper bound, while $\delta$ is the true message delivery time.
Other aspects. Many works such as (Malkhi and Nayak, 2023; Camenisch et al., 2022; Gueta et al., 2019) also focus on the message complexity of the fast path and of the view change. We do describe how BLS signature aggregation (Boneh et al., 2018) can be used for FICC, but leave further advancements to future work. Message complexity and performance do not always go hand in hand in practice (Camenisch et al., 2022) and oftentimes probabilistic broadcast through gossiping is preferred (Protocol Labs, 2023).
3. Model
We consider a network of $n\geq 3f+2p+1$ participants called replicas. Our protocol is safe and live with up to $f$ replicas being Byzantine. Additionally, we guarantee fast termination when up to $p$ replicas are not cooperating. In other words, the parameter $p\in[0,f]$ denotes the maximum number of replicas that are not needed for the fast path to be successful.
In particular, by setting $p=0$, we reach the upper bound on the number of Byzantine replicas permissible (Dwork et al., 1988), i.e. $n\geq 3f+1$. In this case, the fast path will be used only when all replicas, including Byzantine ones, are in pre-agreement on a block and behaving honestly. On the other side of the spectrum, the fast path can be made robust against Byzantine behavior, by setting $p=f$.
We consider the same communication model used in ICC, in which replicas communicate over a partially synchronous network (Dwork et al., 1988). Partial synchrony lies between the definitions of a synchronous network and an asynchronous network. In a synchronous network, there is a known fixed upper bound $\Delta_{bnd}$ (or simply $\Delta$) on the time required for a message to be sent from one party to another. In an asynchronous network no fixed upper bounds $\Delta_{bnd}$ exists. We use $\delta$ to denote the true message delivery time of replicas, i.e., the unknown time it takes for one communication step. In partial synchrony, the network alternates intervals of synchrony, in which the bound $\delta<\Delta_{bnd}$ holds, to ones of asynchrony.
Replicas communicate via all-to-all authenticated links. We assume the existence of a public-key infrastructure (PKI), secure digital signatures, and collision-resistant hash functions. Additionally, a random leader election oracle is assumed to exist.
4. Internet Computer Consensus
The Internet Computer Consensus protocol (ICC) is used to make replicas in the same network agree on a chain of blocks. As the protocol progresses, a tree of blocks is constructed, starting from a genesis block that is at the root of the tree. While each replica may have a different, partial view of the tree, no replicas have a conflicting view of the tree. As the replicas make progress, by finalizing at most one block per tree height, a path of finalized blocks is established. This path constitutes the blockchain. Honest replicas have a consistent view of this path.
Creating the Block-Tree through Notarization. The ICC protocol proceeds in rounds. In the k-th round, one or more blocks are added to the tree at height $k$. In the following, we will use round and height interchangeably. To be added to the tree, a block must be notarized. In ICC, a block $b$ becomes notarized for a replica, once at least $n-f$ notarization shares for block $b$ are received. A notarization share is a signature share of a BLS multi-signature scheme (Boneh et al., 2018) in which a signature can be reconstructed by combining any $n-f$ of the $n$ signature shares, each signed by a different replica. A notarization share sent by a replica $u$ for a block $b$ implies that replica $u$ validated block $b$. As soon as block $b$ is validated by at least $n-f$ replicas, and thus becomes notarized, it can be added to the block-tree.
Block Proposal. In principle, in every round $k$, each replica can propose a block and all notarized blocks make it into the block-tree. Therefore, at any height $k$ there can be up to $n$ blocks. However, this makes it difficult for the replicas to agree on which block should be part of the blockchain, as at most one block per height should be included. To reduce the number of blocks proposed in each round, a random beacon is used to generate a random permutation of the $n$ replicas. The permutation defines a different rank $r\in[0,n-1]$ for each replica for that particular round. The replica with the lowest rank, i.e., the rank-0 replica, is the leader of that round. At the beginning of a round, each replica computes its rank and starts a timer. The leader of that round proposes the block immediately, while the other replicas wait for a time directly proportional to their rank before proposing a block. Let $r$ be the rank of a replica in a given round, the replica delays the proposal of its block by a proposal delay $\Delta_{prop}(r)=2\Delta_{bnd}r$. In round $k$, when deciding which block in the block-tree to extend, the replica considers only notarized blocks at height $k-1$.
Notarization. In round $k$, before sending a notarization share for a block $b$ proposed by a replica $u$ with rank $r^{(u)}$, the replica $v$ receiving the block $b$ waits a notarization delay of $\Delta_{notary}^{(u)}=2\Delta_{bnd}r^{(u)}$ after starting round $k$. Therefore, once replica $v$ receives a block $b$ from replica $u$, it checks if it is time to send a notarization share for a block of rank $r^{(u)}$. If so, it broadcasts the notarization share to all other replicas, otherwise, it waits until $\Delta_{notary}^{(u)}$ has passed. This way, blocks of lower rank should become notarized and added to the block-tree before others. Once a block $b$ in round $k$ becomes notarized for replica $u$, i.e., replica $u$ receives at least $n-f$ notarization shares for $b$ and combines them into a notarization, replica $u$ broadcasts the notarization to other replicas in the subnet, stops sending notarization shares for blocks of round $k$, and starts round $k+1$. Thus, when the leader is honest, the network is in a phase of synchrony and the block proposal and notarization delays are set accordingly, only the block proposed by the rank-0 replica will be added to the tree at height $k$. If the leader is not honest or the network is asynchronous, some other replicas of higher rank may also propose blocks, and also have their blocks added to the tree.
Finalization. As there might be multiple notarized blocks at the same height, we must guarantee that all replicas agree on the same finalized block and that all finalized blocks are part of the same blockchain. Each replica $u$ could assume that after a certain time period, it will have seen all notarized blocks that have been proposed, and if there is just one notarized block at a height $k$, it considers the chain to be agreed upon up to this height. However, this approach does not provide a deterministic finalization guarantee as perhaps the network is in a moment of asynchrony and there are other notarized blocks at height $k$ that the replica $u$ has not received yet. If the replica does not wait long enough, it may violate safety. Thus, on one side blocks need to be agreed upon quickly to reduce the block finalization delay as much as possible, but on the other side, replicas might have to wait a long time to not break consistency. To avoid this trade-off, ICC uses a mechanism similar to notarization to guarantee deterministic finalization. If the replica did not previously broadcast notarization shares for any other block in the same round $k$, the replica will also broadcast another signature share for block $b$, called finalization share. If a replica $u$ receives at least $n-f$ finalization shares for the same block $b$, the replica $u$ can combine them into a finalization. At this point, the replica $u$ is guaranteed that all replicas will eventually include block $b$ in the blockchain. Once a finalization is created for a given block $b$, block $b$ is said to be explicitly finalized and can be output. The finalization is then sent to all the other replicas in the subnet.
The rule for broadcasting finalization shares implies that in some rounds no block might become finalized by the aggregation of at least $n-f$ finalization shares. However, as soon as a block in a later round is explicitly finalized, all its ancestors in the tree, until the last explicitly finalized block, automatically become implicitly finalized. An illustration of the steps necessary before reaching finalization is shown in fig. 3.
Figure 3. The diagram shows the steps necessary to reach the finalization of a block. Black blocks are finalized, dark gray blocks are notarized, while light gray blocks are not. In this example, $n=4$ and $f=1$. Initially, $n-f=3$ replicas send notarization shares (NS) for the rank-0 block. Replicas will not send a NS for higher rank proposals, in this round. As soon as the NS are received, the replica combines them in a notarization (N). Then, the replicas that only sent an NS for the rank-0 block will also send a finalization share (FS) for it. Once the three FS are received, they will be combined into a finalization (F) and both the r-0 block and its parent(s) are finalized.
Lemma 4.1 ().
Let $n\geq 3f+1$, and $b$ be a finalized block at height $k$. ICC guarantees that if block $b^{\prime}$ is notarized at height $k$, then $b=b^{\prime}$.
Proof.
Let a block $b$ at height $k$ be finalized for replica $u$. Replica $u$ must have received at least $n-f$ finalization shares for block $b$, $n-2f$ of which are from honest replicas. Each honest replica sends a finalization share for a notarized block $b$ at height $k$ only if it did not send any other notarization share for other blocks at height $k$. Therefore, other blocks at height $k$ receive at most $2f$ notarization shares. The bound $n>3f$ implies that $2f<n-f$, thus other blocks at height $k$ do not receive enough notarization shares to become notarized. ∎
In other words, if a block $b$ at height $k$ becomes finalized, no other block at height $k$ can be notarized. Since only notarized blocks are extended, lemma 4.1 implies that all blocks at height greater than $k$ will have $b$ as an ancestor and thus, all finalized blocks are part of the same blockchain. This sketches the safety proof of the ICC algorithm.
{remark}
[Finalization Latency] In rounds in which the network is synchronous, finalization of a block can be reached in three times the message delivery time $\delta<\Delta_{bnd}$. In the first round the block proposal, in the second the notarization shares, and in the third the finalization shares are broadcast (see fig. 1).
{remark}
[Timeouts] In the simplest implementation of the ICC protocol, we can assume that the communication delay bound $\Delta_{bnd}$ is an explicit parameter. In practice, instead, the protocol is modified to adaptively adjust to an unknown communication delay bound.
5. Problem Statement
The role of the consensus layer of the Internet Computer is to order blocks containing an arbitrary payload so that all replicas output the payload in the same order. The properties we want the Fast Internet Computer Consensus algorithm to satisfy, are first and foremost the same that Internet Computer Consensus provides (Camenisch et al., 2022):
  • Deadlock Freeness: In each round, the block-tree height increases by 1.
  • Safety: If some honest replica finalizes block $b$ in round $k$, and another honest replica finalizes block $b^{\prime}$ in round $k$, then $b=b^{\prime}$.
  • Liveness: If the network is momentarily synchronous and the leader is honest, then the block proposed by the leader is added to the block-tree and finalized.
Additionally, Fast Internet Computer Consensus satisfies a stronger property:
  • Fast Termination: If the network is momentarily synchronous, the leader is honest, and $n-p$ replicas behave momentarily honestly, then the block proposed by the leader is added to the block-tree and finalized in a single round trip time.
{remark}
Property 5 is also a liveness property. It provides chain-growth even in periods of asynchrony, while Property 5 guarantees that the entire progress is finalized in periods of synchrony. This distinction displays the two modes of operation (Camenisch et al., 2022).
6. Intuition
The goal of the Fast Internet Computer Consensus fast path is to finalize a block as soon as possible while guaranteeing consensus among replicas. In the ICC protocol, a block is finalized at the fastest three communication steps after being proposed. The fast path added in FICC reduces the finalization latency down to two communication steps, as shown in fig. 1.
Definition 6.1 ().
In order to distinguish the two possible finalization scenarios, we call Fast Path finalized (or FP-finalized) the blocks that are explicitly finalized via the FICC fast path, while we call Internet Computer finalized (or IC-finalized) the blocks that are explicitly finalized by receiving at least $n-f$ finalization shares, as in the ICC protocol.
The idea behind the fast path is that if “enough” replicas send their first notarization shares for the same block $b$, i.e., replicas are in pre-agreement, then $b$ can be immediately finalized without waiting for the finalization shares to be sent and received. In other words, replicas fall back to using the finalization through finalization shares only if there are “too many” replicas that do not agree on the first block added to the block-tree at a given height.
Definition 6.2 (Fast Share).
To determine if enough replicas are in pre-agreement, and try to finalize a block via the fast path, replicas broadcast a fast share for the first block they notarize in each round. A block that receives at least $n-p$ fast shares becomes explicitly finalized via the fast path, and fast shares can be combined into a fast finalization.
Safety considerations. As we have sketched in lemma 4.1, the rules of ICC ensure that if a block $b$ at height $k$ becomes IC-finalized, no other block at the same height will be notarized. This guarantees that the tree will not contain any block at height $k$ besides block $b$. Therefore, only block $b$ will be extended. In FICC however, if a block $b$ gets FP-finalized, we cannot guarantee that there will not be other notarized blocks at the same height. If implemented without care, this could induce the finalization of conflicting blocks.
Hence, we introduce a new concept, similar to notarization. Intuitively, we want to guarantee that whenever a block $b$ at height $k$ gets FP-finalized, then no other block at height $k$ can be extended. To this end, each replica marks blocks as fastable when they are safe to be extended.
As we will show, the conditions of blocks being fastable almost never restrict the original ICC algorithm, and only a few additions are sufficient to guarantee the safety of FICC.
7. Fast Internet Computer Consensus
The FICC algorithm is an extension to the Internet Computer Consensus algorithm that enables fast path finalization in periods of synchrony. The protocol provides additional conditions that, if satisfied, guarantee explicit finalization to be achieved as soon as enough fast shares are received. The fast path runs alongside the Internet Computer Consensus protocol and is integrated into existing messages. In case the fast path cannot be used, there is no “switching cost” to revert to the slow path. When there is no pre-agreement, blocks will still be notarized and eventually finalized (either explicitly or implicitly) just as in the ICC protocol.
The following definitions are written from the point of view of a replica $u$:
Definition 7.1 ().
Let B(k) be the set of blocks that have been received in round $k$. Let FS(b) be the set of replicas that have sent a fast share for $b\in B(k)$.
Definition 7.2 ().
At any given time, m(k) is the round-k block for which the most fast shares have been observed:
$m(k)=\operatorname*{arg\,max}_{b\in B(k)}{\left|FS(b)\right|}$
Definition 7.3 ().
At any given time, sum(k) is the number of distinct replicas that have sent a fast share for at least one round-k block:
$sum(k)={\left|\bigcup_{b\in B(k)}FS(b)\right|}$
Using these definitions, we can pin down exactly what blocks are safe to extend for a replica $u$, in the presence of the fast path. We call such a block fastable.
Definition 7.4 (Fastable Blocks).
Finalized blocks are fastable by definition. Additionally, at any point during a round $k$:
  1. (1)
    if ${\left|FS(b)\right|}>f+p$, then $b$ is fastable.
  2. (2)
    if $sum(k)-{\left|FS(m(k))\right|}>f+p$, all current and future blocks of round $k$ are fastable.
An example of the above definitions is presented in fig. 4.
Figure 4. Fast shares (FaS) received for each block are shown in the block-tree. Black blocks are finalized, dark gray blocks are fastable, while light gray blocks are not. Supposing $n=4$, $f=1$, and $p=0$, for round $k$, as $sum(k)=3$, only the first block is fastable. Instead, for round $k+1$, $sum(k)=4$ and as $4-2>1$, all blocks in this round are fastable.
The Fast Internet Computer Consensus algorithm FICC is defined by the following changes to the ICC algorithm presented in section 4:
  1. Restriction 1
    Block proposals can only extend a fastable block. Similarly, notarization shares, fast shares, and finalization shares are only sent for blocks that extend a fastable block.
  2. Restriction 2
    Replicas move to the next round once a fastable block has been notarized.
  3. Addition 1
    When replicas move to the next round, a fastable proof of the notarized block is broadcast.
  4. Addition 2
    When a block is proposed, a fastable proof of the parent block is broadcast alongside the proposed block.
  5. Addition 3
    Alongside the first round-$k$ notarization share, a fast share is broadcast.
  6. Addition 4
    If a block has received $n-p$ fast shares, it is considered (FP-)finalized. Fast shares are combined into a fast finalization and broadcast.
Algorithms 1 and 2 contain the full FICC pseudocode.
{remark}
Fastable proofs (Addition 1) can be implemented naively by aggregating the fast shares for each block (Boneh et al., 2018). In periods of extreme asynchrony, no aggregation might be possible, from which follows a worst-case proof size of $12f+2p+$ 1 shares. In practice, however, we expect only a few different blocks to receive fast shares. The application of more targeted cryptographic primitives to achieve small fastable proofs in the worst case is left to future work.
1:Implements:
2:     Fast Internet Computer Consensus, instance ficc
3:     Code for round-$k$ at replica $u$
4:
5:Uses:
6:     RandomLeaderElectionOracle, instance oracle
7:     BestEffortBraodcast, instance broadcast
8:
9:Parameters:
10:     $\Delta_{prop}$triangleright Proposal Delay
11:     $\Delta_{notary}$triangleright Notarization Delay
12:     $k$triangleright Current round number
13:     $kMax$triangleright Highest round finalized so far
14:     $r_{u}$triangleright Rank assigned by the oracle
15:     $payload$triangleright Set of transactions for this round
16:     $b_{p}$triangleright Notarized and fastable round-($k-1$) block
17:
[…]
19:     $fastSent:=\text{\tt False}$; $\triangleright$ Fast share sent
20:     $proposed:=\text{\tt False}$; $\triangleright$ Proposed a block
21:     $t_{0}:=clock()$; $\triangleright$ Time at the start of the round
22:     $D:=\emptyset$; $\triangleright$ Set of disqualified ranks
23:     $N:=\emptyset$; $\triangleright$ Set of blocks for which a notarization share was sent
24:
25:upon $\neg proposed\;\textbf{and}\;clock()\geq t_{0}+\Delta_{prop}(r_{u})$ do
26:     create a new round-k block:
27:     $b:=(k,u,hash(b_{p}),payload,\sigma_{u})$
28:     $proposed\leftarrow\text{\tt True}$
29:     broadcast $b$, $b_{p}$’s notarization, $b_{p}$’s fastable proof
30:
31:upon exists a valid round-$k$ block $b$ of rank $r$ such that $b\notin N,\;\textbf{and}\;r\notin D\;\textbf{and}\;clock\geq t_{0}+\Delta_{notary}(r)\;\textbf{and}\;\nexists\textit{ valid round-$k$ block $b^{\prime}$ of rank $r^{\prime}$}\in[r]\setminus D$ do
32:     if $r_{u}\neq r$ then
33:         broadcast $b$, $b_{p}$’s notarization, $b_{p}$’s fastable proof
34:     end if
35:     if some block in $N$ has rank $r$ then:
36:         $D\leftarrow D\cup\{r\}$;
37:     else
38:         $N\leftarrow N\cup\{b\}$
39:         if $\neg fastSent$ then:
40:              broadcast fast share for $b$;
41:              $fastSent\leftarrow\text{\tt True}$;
42:         end if
43:         broadcast notarization share for $b$
44:     end if
45:
Algorithm 1 FICC (Part 1)
[…]
2:     combine notarization shares into a notarization for $b$
3:
[…]
5:     combine fast shares into a fastable proof
6:     broadcast notarization and fastable proof for $b$
7:     if $N\subseteq\{b\}$ then
8:         broadcast finalization share for $b$
9:     end if
10:     $k\leftarrow k+1$triangleright Move to next round
11:
[…]
13:     combine (fast) finalization shares into a (fast)
finalization for $b$, if necessary
14:     broadcast (fast) finalization for $b$
15:     output payloads of the last $k-kMax$ blocks in the chain ending at $b$
16:     $kMax\leftarrow k$
17:
18:procedure valid($b$) is
19:     return $b$ extends a notarized and fastable round-$(k-1)$ block $b_{p}$ and $b$ is signed correctly
20:
Algorithm 2 FICC (Part 2)
8. Protocol Analysis
8.1. Deadlock Freeness
The ICC protocol guarantees deadlock freeness as each honest replica receives at least one notarized block in each round. Therefore, each honest replica can start the following round and the block-tree keeps growing. However, in FICC  a block that was notarized is not guaranteed to be extended as, due to Restriction 1, in order for this to happen, the block must also be fastable.
The following theorem proves that regardless, deadlock freeness is not impacted. Intuitively, Restriction 2 makes sure that replicas only move to a higher round once a block can be extended, and Addition 1 guarantees that each replica does so eventually.
Theorem 8.1 ().
FICC satisfies deadlock freeness.
Proof.
We prove that the block-tree of each honest replica keeps growing in height by induction on the block-tree height $k$. Specifically, we show that for each round $k$, at least one notarized and fastable block will be added to each honest replica’s block-tree.
For the base case, note that all honest replicas agree on the root of the block-tree, the genesis block, by definition. This genesis block is notarized and finalized by definition too, which in turn makes it fastable by definition 7.4. Thus, in round 0, a notarized and fastable block is added to each node’s block-tree.
Now let’s assume that a notarized and fastable block exists in round $k$. We prove that each replica observes a notarized and fastable block in round $k+1$. First, note that all honest replicas will enter round $k+1$, by executing lines 49 to 55, as by the induction hypothesis they all have a block that can be extended.
Each honest replica will receive at least one round-(k+1) block, together with a notarization and fastable proof for the extended parent block (Addition 2). Thus, each honest replica will send a fast share for one block (Addition 3). Eventually, each replica is thus guaranteed to observe $n-f$ fast shares. Note that at the latest at this point, at least one block will be fastable, since either (i) one block receives more than $f+p$ fast shares, or (ii) other blocks receive more than $n-f-(f+p)>f+p$ fast shares. As no honest replica moves to a higher round without a fastable block being notarized (Restriction 2), and replicas keep notarizing blocks, eventually, one fastable block will be notarized. Honest nodes moving from one round to the next broadcast all the information needed for other honest nodes also to receive a notarized and fastable block (Addition 1).
{remark}
Any scenario in which notarized blocks for round-k exist before any of them are fastable can only occur due to message reordering. Hence Restriction 1 and Restriction 2 cannot cause latency concessions in the slow path if the communication channel precludes message reordering.
8.2. Safety
We guarantee that all honest replicas agree on the chain of finalized blocks. Suppose an honest replica finalizes block $b$ at height $k$ and let $pt$ be the path in the block-tree from the genesis to $b$. Suppose another honest replica finalizes block $b^{\prime}$ at height $k^{\prime}$, where $k^{\prime}\geq k$, and let $pt^{\prime}$ be the path in the block-tree from genesis to $b^{\prime}$.
Requiring that all honest replicas agree on the finalized blocks is equivalent to requiring that $pt$ be a prefix of $pt^{\prime}$. This is guaranteed as long as, if a block is finalized (either IC-finalized or FP-finalized), any other block at the same height can be neither finalized nor extended. Thus, if a block at height $k$ is IC-finalized, there must not be another block at height $k$ which is notarized. Similarly, if a block at height $k$ is FP-finalized, there must not be another block at height $k$ which is fastable.
Lemma 8.2 ().
Let $n\geq 3f+2p+1$ , in any round, if a block $b$ is IC-finalized, then FICC guarantees that no block $b^{\prime}$, $b^{\prime}\neq b$ is notarized for any replica.
Proof.
As no changes have been made to FICC in terms of notarization, this statement follows directly from lemma 4.1. ∎
Therefore, any other block at height $k$ will be neither extended nor finalized (as both cases require it to be at least notarized). Similarly, let $b$ be a block at height $k$ which is FP-finalized, we prove that $b$ is the only fastable block at height $k$.
Lemma 8.3 ().
Let $n\geq 3f+2p+1$ , in any round, if a block $b$ is FP-finalized, then FICC guarantees that no block $b^{\prime}$, $b^{\prime}\neq b$ is fastable for any replica.
Proof.
Assume towards contradiction that there exists an FP-finalized block $b$ and a fastable block $b^{\prime}$ at height $k$. By the definition of FP-finalization a replica observed $|FS(b)|\geq n-p\geq 3f+p+1$ fast shares for $b$. Thus, all blocks different from $b$ have received at most $f+p$ fast shares in total. In particular $|FS(b^{\prime})|\leq f+p$. This means that condition (1) of definition 7.4 is not met and hence condition (2) must hold for $b^{\prime}$ to be fastable: $sum(k)-{\left|FS(m(k))\right|}>f+p$.
  • Let’s assume $m(k)=b$, then $sum(k)>f+p+|FS(b)|$, which implies that more than $p$ replicas have not sent a fast share for $b$, thus reaching a contradiction.
  • Instead, let’s assume $m(k)=b^{\prime}$, $b^{\prime}\neq b$. Then ${\left|FS(b)\right|}\leq{\left|FS(b^{\prime})\right|}$ by the definition of $m(k)$. Thus $sum(k)>f+p+|FS(b^{\prime})|$ implies $sum(k)>f+p+|FS(b)|$, which is a contradiction since there can be at most $p$ honest replicas not sending a fast share for $b$.
Figure 5. The rank-1 (r-1) block is fastable (Fa) for replica $u$, while the rank-0 (r-0) block is FP-finalized for replica $v$. Lemma 8.3 proves that this situation cannot arise.
Theorem 8.4 ().
FICC satisfies safety.
Proof.
We start by noting that FP-finalization implies notarization, hence IC- and FP-finalization cannot happen on different blocks of the same round. Lemma 8.2 proves that, if a block is IC-finalized, then there is no other notarized block at the same height. Likewise, lemma 8.3 proves that, if a block is FP-finalized, then there is no other fastable block at the same height.
In both cases, since only notarized and fastable blocks are extended, if a block at height $k$ is finalized, no other blocks at height $k$ will be extended and no fast, notarization, or finalization shares will be sent for these blocks (Restriction 1). This guarantees that whenever a block might get finalized all replicas only extend this block. Thus finalized blocks are always part of the same blockchain. ∎
8.3. Liveness
We define liveness such that whenever the network is synchronous for a “long enough” interval, if the leader is honest, only the leader’s block will be added to the block-tree for the corresponding round and this block will be finalized by all honest replicas.
Theorem 8.5 ().
FICC satisfies liveness.
Proof.
Let $T$ be the time when the first honest replica $u$ enters round $k$. Suppose that the rank-0 replica $q$ of round $k$ is honest. Moreover, suppose that the network is $\delta$-synchronous between time $T$ and $T+2\delta+\Delta_{prop}(0)$. Furthermore, assume $\delta\leq\Delta_{bnd}$, which implies that $\Delta_{notary}(1)\geq 2\delta+\Delta_{prop}(0)$, with the functions $\Delta_{notary}$ and $\Delta_{prop}$ as defined in section 4.
We want to prove that under these assumptions, eventually, each honest replica will see replica $q$’s proposed block as finalized.
Before the honest replica $u$ enters round $k$ at time $T$, the replica $u$ notarized a fastable block in rounds $11,\dots,k-$ 1 and broadcast the notarization and fastable proof (Addition 1). By the synchrony assumption, the other replicas, $q$ included, receive the notarization and fastable proof for rounds $11,\dots,k-$ 1 by time $T+\delta$, and enter round $k$. Thus, replica $q$ broadcasts its rank-0 proposal by time $T+\delta+\Delta_{prop}(0)$, and other replicas will receive it by $T+2\delta+\Delta_{prop}(0)$. By assumption no other honest replica enters round $k$ before time $T$ and as $\Delta_{notary}(1)\geq 2\delta+\Delta_{prop}(0)$, every honest replica will broadcast a fast share and a notarization share for the rank-0 block (Addition 3). Therefore, all honest replicas eventually receive at least $n-f$ fast shares for the same rank-0 block. When $p=f$, these fast shares are enough to FP-finalize the block proposed by replica $q$ (Addition 4). When $p<f$, ${\lfloor\frac{n+f}{2}\rfloor}+1\leq n-f$ notarization shares are sufficient to notarize the block instead. As argued previously, honest replicas will not have notarized another block and thus will broadcast a finalization share. Eventually, after an additional communication round honest replicas will receive ${\lfloor\frac{n+f}{2}\rfloor}+1\leq n-f$ finalization shares, which leads to IC-finalization. ∎
8.4. Fast Termination
Theorem 8.6 ().
FICC satisfies fast termination.
Proof.
With the additional assumption that $n-p$ replicas behave honestly during a phase of synchrony, we can guarantee $n-p$ fast shares for the rank-0 block reaching every replica. Thus, after a single round-trip time, each replica will be able to FP-finalize the rank-0 block. ∎
9. Evaluation
9.1. Methodology
As mentioned in section 8.1, a scenario with a different block proposal timing between FICC and ICC can only be crafted by reordering some of the messages exchanged between replicas. Therefore, in all experimental scenarios that address normal network conditions or enforce message order (e.g. use of TCP), FICC and ICC will exhibit identical behavior and timing with respect to dissemination and notarization of blocks. While FICC will of course behave differently to finalize received blocks, throughput is determined entirely by block creation interval, which remains unchanged.
We conclude that the throughput performance is unchanged in FICC  and bigger block payloads have the same effects on throughput and latency in FICC and ICC. To isolate the effects of the changes between the two protocols, we focus the experiments on the finalization latency of small blocks.
Application subnets of the Internet Computer feature 13 replicas (Dfinity, 2023a). To test the potential of FICC as a drop-in replacement for ICC, and since $n=16$ is optimal for both $f=5$, $p=0$ and $f=3$, $p=3$ experimental scenarios, we perform experiments with up to 16 replicas around the globe (see fig. 6). Replicas run on AWS t3.medium EC2 instances (with 2vCPUs, 4 GB of memory, running Ubuntu 22.04).
Figure 6. The coordinates of the 16 deployed EC2 instances forming a global network are marked by black circles (section 9.3). The red triangles mark 13 clustered replicas and 3 eccentric replicas (section 9.4), while the blue squares show the coordinates of 4 distant replicas (section 9.5).
We wish to demonstrate that (i) in any network topology FICC is performing consistently as well as or better than ICC, while (ii) outperforming ICC in evenly geo-distributed settings. Finally, we want to show (iii) the positive effect of setting $p>0$, allowing for latency improvements even in the face of bottlenecks such as eccentric nodes.
As the goal of our evaluation is to measure the precise impact of our changes, we define latency as the proposal finalization time, both at the proposer (rank-0 replica) and at the other replicas. While the proposer finalization time can be measured precisely using the proposer’s system clock, other replicas have no information about when a round was started. We circumvent this inherent limitation, by relying on clock accuracy across AWS datacenters, which is handled by periodic synchronization to an NTP server (Amazon, 2023). This allows us to use the block timestamp as a round’s starting time.
As required by the protocol, we set the proposal ($\Delta_{prop}$) and notarization ($\Delta_{notary}$) delays to be larger than the message delay experienced without network disruptions, such that our experimental results correspond to regular network conditions and only one block is proposed per round. Each experiment is run for 100 seconds, which is shown to be sufficient, as the measurements show remarkable periodicity (e.g., see fig. 7(b)).
9.2. Implementation
We implement FICC in Rust by adapting the existing implementation of the Internet Computer Protocol (Dfinity, 2023b) and open-source our code (Massimo Albarello, 2023). To test performance we strip the original codebase of all but its consensus layer and use libp2p’s flood-sub implementation (Protocol Labs, 2023) as the network layer. For clarity and simplicity, we also replace random beacon leader election with round-robin rotation.
9.3. Global Network
(a) Proposal finalization times across all replicas, for ICC and two FICC variants.
(b) FICC with f=5, p=0. Proposal finalization time at replica in Tokyo (ap-notheast-1). Own proposals are marked in red.
(c) FICC with f=5, p=0. Proposal finalization time at respective proposer.
Figure 7. Evaluation results of a network consisting of 16 globally distributed replicas.
We simulate a worldwide deployment akin to an IC subnet by distributing 16 replicas as shown by the black circles in fig. 6.
In this setting the average proposal finalization time measured for the ICC implementation was 845ms. By running FICC with $f=5,p=0$, an average proposal finalization time of 705ms is measured instead, while FICC with $f=3,p=3$ further reduces it to 638ms (fig. 6(a)).
Setting $f=5$ and $p=0$ leads to better results compared to ICC, as most replicas can FP-finalize blocks. In fig. 6(b), the proposal finalization time is shown from the point of view of the replica situated in Tokyo. Interestingly, one can observe that the fast path is almost always taken, except for its own proposals (marked by red dotted lines). Intuitively, this can be explained by the fact that in a highly connected network topology such as this one, fast shares from replicas situated between the proposer and recipient arrive just after the block proposal. (We hypothesize that this would not happen in e.g. a star network topology.)
Figure 6(c) shows that for some rounds IC-finalization (i.e., receiving two quorums of $n-f$ shares each) was still faster than waiting for a single quorum of $n-p$ fast shares. This is to be expected, especially for proposers on the periphery. Note also that fig. 6(a) shows a smaller variance in latency for ICC. This can be explained as ICC only waits for the $n-f=11$ fastest responses, while FICC requires $n-p$ shares to utilize the fast path (16 and 13 shares respectively). Thus, a remote node is expected to benefit less from the fast path. On the other hand, a node close to the center of the network will benefit more from the reduced number of communication steps as shown in fig. 6(a).
(a) Proposal finalization times across all replicas, for ICC and two FICC variants.
(b) FICC with f=5, p=0. Proposal finalization time as measured by the respective proposer.
(c) FICC with f=3, p=3. Proposal finalization time as measured by the respective proposer.
(d) FICC with f=5, p=0. Proposal finalization time across replicas.
(e) FICC with f=3, p=3. Proposal finalization time across replicas.
Figure 8. Evaluation results of a network consisting of 16 replicas, with 3 eccentric replicas.
9.4. Global Network with Eccentric Replicas
In this experiment, we demonstrate the effect of outliers (i.e., eccentric nodes) on the performance of FICC. As in section 9.3, 16 nodes are globally distributed, however, all except for three nodes are clustered in North America and Europe (see the red triangles in fig. 6). As fig. 7(a) shows, ICC averages a proposal finalization time of 853ms, which FICC with $p=0$ only marginally improves on (avg. finalization time of 844ms). This is to be expected, as the clustered nodes would need to wait for all three outliers to finalize a block through the fast path. Instead, they can IC-finalize blocks amongst themselves, without input needed from the outliers.
When the parameter $p$ is raised to $p=3$, the closely connected replicas can finalize blocks through the fast path by themselves, which explains the 20.4% drop in latency they experience in this scenario (679ms finalization time). Figure 7(d) and fig. 7(e) confirm that three replicas experience high latency. Even when $p=3$, their latency isn’t improved, however, the large cluster vastly benefits, as they now can finalize through the fast path, without having to wait for finalization shares from the eccentric replicas.
9.5. Global Network with Few Distant Replicas
Figure 9. Proposal finalization times across all replicas, for ICC and FICC respectively, obtained by benchmarking a network of 4 replicas for 100 seconds.
In this experiment, we measure the potential for FICC improvements under the adversity of an almost maximal message delay in a small network. Four replicas are distributed as shown by the blue squares in fig. 6. Due to the network’s size $p=0$ is forced. As fig. 9 shows, ICC averages a proposal finalization time of 1.30 seconds, which FICC manages to improve by 15.1% to 1.10 seconds. In absolute terms, this is a much higher latency than the previous experiments and shows that geolocation plays a major role when comparing algorithms in practice.
10. Conclusion
The Internet Computer Consensus protocol family borrows techniques from blockchains in order to improve the complexity of classical BFT protocols. The proposed Fast Internet Computer Consensus improves upon the original algorithm by providing a simultaneous dual mode, without incurring extra cost, thus closing the gap between state-of-the-art SMR protocols and classical fast consensus literature.
Concretely, we show that it is possible to achieve optimally fast proposal finalization time for a rotating leader SMR protocol. We experimentally demonstrate the effectiveness of the fast path by showing latency improvements compared to ICC. Ultimately, the effectiveness of the fast path depends largely on the network topology, as well as the parameter $p\in[0,f]$. If $p=0$, then a single corrupt replica can prevent fast path finalization by not sending its fast share. In the context of permissioned blockchains though, this setting is compelling, since it is possible to detect adversarial behavior and remove such misbehaving replicas (Buterin and Griffith, 2017; Camenisch et al., 2022). Moreover, in case a fast path robust to adversarial behavior or straggling replicas is desired, $p$ can be set arbitrarily close to $f$. This allows to take advantage of FICC even in the setting of persistent adversity and unfavorable network conditions.